OpenStudy (anonymous):

Simplify: 6m + 7n + 5m – 3n

OpenStudy (texaschic101):

you can only add like terms 6m + 7n + 5m - 3n 6m + 5m = 11m 7n - 3n = 4n giving us : 11m + 4n

OpenStudy (anonymous):

Which are the like terms in 2y3 – 4y2 + y + y3?

OpenStudy (anonymous):

2y^3-4y^2+y+y^3?

OpenStudy (texaschic101):

like terms : 2x and x, or 3x^2 and 4x^2, or x^3 and 2x^3 so your problem....2y^3 - 4y^2 + y + y^3 the only like terms in this problem is the y^3's answer is : 3y^3 - 4y^2 + y (because 2y^3 + y^3 = 3y^3...there are no other like terms)

OpenStudy (anonymous):

What is the value of the expression 7–3?

OpenStudy (anonymous):

7^-3